If the perimeter is 86,and the length is 25, then what is the width? I've tried a rule that said the width is four times less than the length, but that did not work for this problem because the perimeter added up to 92 instead of 86 inches.
Answer by macston(5194)

.
W=width; L=length=25; P=perimeter=86
P=2(L+W) This is one formula to find perimeter.
P/2=L+W
(P/2)-L=W Plug in values and solve.
(86/2)-25=W
43-25=W
18=W
CHECK:
P=2(L+W)
86=2(25+18)
86=2(43)
86=86